OpenStudy (kj4uts):

The law of cosines reduces to the Pythagorean theorem for right triangles. True or false? Please explain. Thank you!

OpenStudy (kj4uts):

Oh ok I am just confused because I found places that say the answer is false so im not sure what the answer is?

OpenStudy (kj4uts):

But I guess this person says true so I guess there is more true than false? https://answers.yahoo.com/question/index?qid=20140520093037AAr7yjI

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

\[c^2=a^2+b^2−2ab~ \cos(90^o)\] so yes it would be true :)

OpenStudy (kj4uts):

Ok thank you both :)

pooja195 (pooja195):

no problem ^_^

OpenStudy (anonymous):

pooja is the one that did the work I didn't do anything. thanks @pooja195 !